# The viscoelastic rheology of transient diffusion creep

### v1

@@ 1 Introduction
The aim of the present manuscript is to provide a simple grain-scale model of rock viscoelasticity that can be used as a basis for interpreting and extrapolating laboratory experiments. The model describes the effective viscoelastic behaviour which arises from the grain-scale diffusion of matter by physical mechanisms referred to as diffusionally-assisted or diffusionally-accommodated grain-boundary sliding [1, 2]; mechanisms also referred to as transient diffusion creep. There is a long history of models of transient diffusion creep, with early pioneering work by Lifshits, Raj and Ashby [3, 4, 5]. The present work builds most closely on the more recent study by Lee et al. [6], who developed a two-dimensional finite element model of a bicrystal consisting of two elastic grains with a grain boundary between. The grain boundary allows for both sliding and diffusion of matter. Here, we model the same physics as Lee et al. [6] but in a different geometry, considering an infinite tiling of hexagonal grains in two-dimensions and a tessellation of tetrakaidecahedral grains in three-dimensions, in turn building on models for steady-state diffusion creep described in [7, 8]. A homogenisation approach is used to determine the effective upscaled viscoelastic tensors which describe the macroscale rheology arising from the grain-scale physics.
@@ 3 Hexagonal grains
Let us consider the end-member behaviour at zero and infinite frequency. At zero frequency (infinite period) the diffusion process reaches a steady-state. This steady-state deformation is steady-state diffusion creep (Coble creep [9]) and the material behaves as a Newtonian viscous fluid. The steady-state shear viscosity $\eta$ for an array of hexagonal grains has been calculated by many authors [e.g. 10, 11, 7] and is
@@
Figure 1: Properties of the model for hexagonal grains as a function of Poisson’s ratio $\nu$. a) Scaled unrelaxed shear modulus $G_{0}/\mu$ and scaled steady-state shear modulus $G_{ss}/\mu$. b) Relaxation strength $\Delta$ associated with grain boundary diffusion, $\Delta=(G_{0}/G_{ss})-1$. c) Ratio of the Andrade time $\tau_{A}$ to the Maxwell time $\tau_{M}$. The data in these plots can be well fit by simple rational functions of $\nu$ (see Appendix J). The plots show the behaviour for the full allowable range of Poisson’s ratio for linear elasticity, $-1\leq\nu\leq\tfrac{1}{2}$; a typical value for rock is $\nu\sim 0.3$.
@@
Figure 2: A summary of the viscoelastic response for hexagonal grains with a Poisson’s ratio $\nu=0.3$. a) $|G^{*}|/G_{0}$, the ratio of the absolute shear modulus to the unrelaxed modulus as a function of dimensionless frequency $\omega\tau_{M}$. b) $Q^{-1}$, inverse quality factor (attenuation) as a function of dimensionless frequency. c) Scaled compliances $(G_{0}J_{1}-1)/\Delta$ and $(G_{0}J_{2}-1/\omega\tau_{M})/\Delta$ as a function of dimensionless frequency. d) Cole-Cole plot of the scaled compliances against each other. In all plots solid lines show the finite element solutions, dashed lines an extended Burgers model fit, dotted lines an Andrade model fit, dash-dotted lines a Maxwell model fit
Figure 2 shows plots which summarise the viscoelastic behaviour of the model for hexagonal grains. The top two panels show quantities commonly plotted in experimental studies of forced oscillation; the bottom two panels show scaled versions of the compliances which illustrate the nature of the relaxation. Figure 2 also shows fits to the finite element results using some simple viscoelastic models: the Maxwell model, the Andrade model and the extended Burgers Model (see Appendix H for a description of these models).
The behaviour of the model is best understood by considering its extremes. As remarked earlier, the behaviour at zero frequency is that of a viscous fluid. The behaviour at low frequencies $\omega\tau_{M}\ll 1$ more generally (the diffusionally-accommodated regime) is that of a Maxwell model with a dashpot of viscosity $\eta$ in series with a spring of compliance $J_{ss}$ or modulus $G_{ss}=1/J_{ss}$. $\eta$ is the steady-state viscosity and $G_{ss}$ is the steady-state modulus. The steady-state compliance $J_{ss}$ is the limiting value of the storage compliance: $J_{1}\rightarrow J_{ss}$ as $\omega\tau_{M}\rightarrow 0$. The relaxation strength $\Delta$ is the fractional increase in compliance due to the relaxation of stresses by grain boundary diffusion, given by
@@
At high frequencies (the diffusionally-assisted regime), $\omega\tau_{M}\gg 1$, the behaviour is that of an Andrade model, with power-law variation of attenuation as $Q^{-1}\propto\left(\omega\tau_{M}\right)^{-\alpha}$ for a particular exponent $\alpha$. The Andrade exponent $\alpha$ can be calculated explicitly using singularity analysis [13, 14, 15] following a scaling argument due to Lee et al. [6]. In the absence of grain boundary diffusion, the pure elastic problem with freely-sliding grain-boundaries has singularities in the stress at the triple junctions where three grains meet, i.e. the normal stresses go as $\sigma_{nn}\propto r^{-\lambda}$ where $r$ is the distance from the triple junction. The analysis of Picu and Gupta [13] shows that for a triple junction with equal angles of 120° and freely sliding grain boundaries the exponent $\lambda=0.4491862$.
In the presence of diffusion, the singularity in the stress is removed. The length scale over which diffusion acts depends on the frequency of the forcing and scales with
@@
The other parameter which specifies the Andrade model is the Andrade time $\tau_{A}$ (appendix H.4). Unlike the exponent $\alpha$, this time cannot be determined purely by examining the singular behaviour at the triple junctions, but is also sensitive to the overall grain geometry and Poisson’s ratio. Figure 1c plots the Andrade time calculated from the finite element analysis as a function of Poisson’s ratio, and shows that the Andrade time is broadly similar to the Maxwell time, increasing from $\tau_{A}/\tau_{M}=0.36$ to 2.13. For a typical rock value of $\nu=0.3$, $\tau_{A}/\tau_{M}=1.315$.
The Maxwell model and the Andrade model characterise the behaviour of the finite element results at low and high dimensionless frequency respectively. Also shown in Figure 2 is another model fit to the finite element data using an extended Burgers Model (appendix H.5). This model is similar to the Andrade model in having a relaxation spectrum with a power-law distribution of relaxation times, but has a long-period cut-off to the relaxation spectrum given by a parameter $\tau_{H}$ which can be related to $\tau_{A}$ and $\Delta$ by $\tau_{H}=\tau_{A}\left(\Gamma(1-\alpha)\Delta\right)^{1/\alpha}$, where $\Gamma(z)$ is the Gamma function. The extended Burgers model has the same asymptotic behaviour as the finite element model at both low and high frequencies. As can be seen in Figure 2, the extended Burgers model provides an excellent fit to the finite element results, with only subtle differences around $\omega\tau_{M}=20$ to be seen in panels a and b which show absolute shear modulus and attenuation. The differences are somewhat more noticeable in the plots of scaled compliance in panels c and d, but even in these plots the differences are slight. Thus an effective parametrisation of the finite element results is to use an extended Burgers model with the appropriate values of $G_{0}$, $\tau_{M}$, $\Delta$ (Figure 1b), $\alpha=0.3672092$ and $\tau_{A}/\tau_{M}$ (Figure 1c).
@@ 5 Discussion
A more significant difference between the present study and that of Lee et al. [6] is the values of $\tau_{A}/\tau_{M}$ which are significantly larger for the bicrystal models of [6] than the models here; that is to say the bicrystal models of [6] show significantly less attenuation for the same dimensionless frequency (the grey lines are lower than the black lines on Figure 6b). This difference must arise from the overall differences in grain geometry, where it should be noted that the bicrystal model has two lengthscales which define it rather than just one for model here: there is one lengthscale associated with the variations of the interface, and another associated with the width of the bicrystal. Lee et al. [6] present results where the width is 5 times that of the scale of variations on the interface; it is expected that the ratio of these two lengthscales plays an important role in determining the amount of attenuation, as dissipation only takes place on the boundary but elastic energy is stored throughout the grains.
@@
Figure 7: Andrade exponent $\alpha$ as a function of junction angle $\theta$ for a triple junction with angles $\theta$, $2180\text{\textdegree}-\theta/$ 2 and $2180\text{\textdegree}-\theta/$ 2. Dotted lines indicate the junction angles for the tetrakaidecahedron ($\theta=109.47\text{\textdegree}$, $\alpha=\tfrac{1}{3}$) and the hexagon ($\theta=120\text{\textdegree}$, $\alpha=0.3672092$).
It is not clear why some laboratory studies show lower Andrade exponents than the models here. In the models, the Andrade exponent is purely controlled by the triple junction geometries, and significant departures from the symmetric 120°-120°-120° junction angles are needed to produce Andrade exponents of 0.2. This is illustrated in Figure 7, which plots the Andrade exponent expected for asymmetric junctions with one angle of $\theta$ and two identical angles of $2180\text{\textdegree}-\theta/$ 2. An exponent of 0.2 could be produced by having a $60\text{\textdegree}-150\text{\textdegree}-150\text{\textdegree}$ junction, but this seems far from what is observed in the microstructures. There may be some other source of anisotropy or geometrical feature that influences the exponent in some experiments, or there may be another dissipation mechanism producing attenuation that is mapped by the fitting to produce a lower effective Andrade exponent.
@@
The relevance and importance of elastically-accommodated grain-boundary sliding in experimental studies is debated. Many studies fit observations by including a high frequency attenuation peak, which some ascribe to elastically-accommodated grain boundary sliding (e.g. [21]) and others do not (e.g. [17, 24]). As noted by Takei [17] (their equation (22)) and Jackson [26] (their equation (13)), a simple model of grain boundary viscosity by Ashby [27] would suggest that the ratio $\tau_{e}/\tau_{M}$ should be around $b^{2}/d^{2}$ where $b\sim 10^{-10}$ m is the atomic size and $d$ is the grain size. For 1 $\mu$m grains, $\tau_{e}/\tau_{M}\sim 10^{-8}$; for 10 $\mu$m grains, $\tau_{e}/\tau_{M}\sim 10^{-10}$; and for 1 mm grains $\tau_{e}/\tau_{M}\sim 10^{-14}$. Thus any peak from elastically-accommodated grain boundary sliding needs data at very high frequencies relative to the Maxwell time to be resolved. The peak that Qu et al. [21] report has $\tau_{e}/\tau_{M}=7\times 10^{-6}$ and a relaxation strength $\Delta_{e}=0.012$. This is a significantly lower relaxation strength than the simple models here predict. The differences in relaxation strength might be attributed due to geometrical factors not modelled here, such as heterogeneity in grain size in real materials [25, 28].
@@
Understanding the viscoelastic behaviour of rocks at high homologous temperatures is of particular relevance to Earth sciences when interpreting the significant variations in seismic wavespeed, as well as increased attenuation, that occur near the lithosphere-asthenosphere boundary [e.g. 29, 23]. The models here do not produce enough attenuation to explain what is seen in seismology. For example, with a steady-state asthenospheric shear viscosity $\eta=2\times 10^{19}~{}$a s and an intrinsic shear modulus $\mu=70$ GPa, the Maxwell time $\tau_{M}=3\times 10^{8}$ s, i.e. around 10 years. For a surface wave with a period around $T=50~{}$s, the dimensionless frequency is $\omega\tau_{M}=2\pi\tau_{M}/T=4\times 10^{7}$, for which the models yield $Q^{-1}\sim 10^{-3}$ (Figure 6), significantly less than the asthenospheric $Q^{-1}\sim 10^{-2}$ estimated by Ma et al. [29] and given in PREM [30]. Previous work in extrapolating laboratory experiments by Jackson and Faul [19] suggested that diffusionally-accommodated grain boundary sliding can account for the $Q^{-1}\sim 10^{-2}$ level of attenuation at seismic frequencies. This extrapolation was based on a proposed variation of Maxwell time as linear with grain size, $\tau_{M}\propto d$, rather than $\tau_{M}\propto d^{3}$ as expected from theory for grain-boundary diffusion. More recent experiments in the same laboratory [31] have since observed a much stronger grain size dependence with $\tau_{M}\propto d^{3.17}$, closer to the behaviour predicted by theory, and for which extrapolation does not yield sufficient attenuation at seismic frequencies. Thus additional mechanisms of dissipation, such as those associated with premelting [18, 24], impurities [32], elastically-accommodated grain boundary sliding [33], or dislocations [34] are needed to explain the seismic observations.
@@
There a number of natural ways the present modelling work can be extended. One natural extension is to consider diffusion within the interior of grains rather than along the grain boundaries. An example of such a calculation for hexagons is given in Appendix L, and the most notable result is that the relationship between the Andrade exponent and the stress singularity exponent changes to $\alpha=1-\lambda$, such that $\alpha=1/2$ is expected for tetrakaidecahedral grains. Calculations could be performed using a synthetic rock texture which has a range of grain sizes, e.g. using Neper [35]. The expectation would be that the Andrade exponent would remain determined by the triple junction angles, but that the effective $\tau_{A}/\tau_{M}$ and $\Delta$ parameters may change. All the calculations here are for materials in solid-state; solid both in the grain interiors and on the grain boundaries. The effect of melt and melt geometry [16] could be explored, as was done for the case of steady-state diffusion creep viscosity in [7]. The presence of melt allows for a steady-state bulk viscosity of the material, and so calculations with melt would allow for both shear and bulk attenuation phenomenon to be explored. However, perhaps most interesting would be to consider the temperature dependent properties of the grain boundaries and the triple lines and explore the effect of premelting and sub-solidus mechanisms of attenuation more generally.

### v2

@@ 1 Introduction
The aim of the present manuscript is to provide a simple grain-scale model of rock viscoelasticity that can be used as a basis for interpreting and extrapolating laboratory experiments. The model describes the effective viscoelastic behaviour which arises from the grain-scale diffusion of matter by physical mechanisms referred to as diffusionally-assisted or diffusionally-accommodated grain-boundary sliding [1, 2, 3]; mechanisms also referred to as transient diffusion creep. There is a long history of models of transient diffusion creep, with early pioneering work by Lifshits, Raj and Ashby [4, 5, 6]. The present work builds most closely on the more recent study by Lee et al. [7], who developed a two-dimensional finite element model of a bicrystal consisting of two elastic grains with a grain boundary between. The grain boundary allows for both sliding and diffusion of matter. Here, we model the same physics as Lee et al. [7] but in a different geometry, considering an infinite tiling of hexagonal grains in two-dimensions and a tessellation of tetrakaidecahedral grains in three-dimensions, in turn building on models for steady-state diffusion creep described in [8, 9]. A homogenisation approach is used to determine the effective upscaled viscoelastic tensors which describe the macroscale rheology arising from the grain-scale physics.
@@ 3 Hexagonal grains
Let us consider the end-member behaviour at zero and infinite frequency. At zero frequency (infinite period) the diffusion process reaches a steady-state. This steady-state deformation is steady-state diffusion creep (Coble creep [10]) and the material behaves as a Newtonian viscous fluid. The steady-state shear viscosity $\eta$ for an array of hexagonal grains has been calculated by many authors [e.g. 11, 12, 8] and is
@@
Figure 1: Properties of the model for hexagonal grains as a function of Poisson’s ratio $\nu$. a) Scaled unrelaxed shear modulus $G_{0}/\mu$ and scaled steady-state shear modulus $G_{ss}/\mu$. b) Relaxation strength $\Delta$ associated with grain boundary diffusion, $\Delta=(G_{0}/G_{ss})-1$. c) Ratio of the Andrade time $\tau_{A}$ to the Maxwell time $\tau_{M}$. The data in these plots can be well fit by simple rational functions of $\nu$ (see Appendix J). The plots show the behaviour for the full allowable range of Poisson’s ratio for linear elasticity, $1\leq\nu\leq\tfrac{1}{2}$; a typical value for rock is $\nu\sim 0.3$.
@@
Figure 2: A summary of the viscoelastic response for hexagonal grains with a Poisson’s ratio $\nu=0.3$. a) $|G^{*}|/G_{0}$, the ratio of the absolute shear modulus to the unrelaxed modulus as a function of dimensionless frequency $\omega\tau_{M}$. b) $Q^{-1}$, inverse quality factor (attenuation) as a function of dimensionless frequency. c) Scaled compliances $(G_{0}J_{1}-1)/\Delta$ and $(G_{0}J_{2}-1/\omega\tau_{M})/\Delta$ as a function of dimensionless frequency. d) Cole-Cole plot of the scaled compliances against each other. In all plots solid lines show the finite element solutions, dashed lines an extended Burgers model fit, dotted lines an Andrade model fit, dash-dotted lines a Maxwell model fit.
Figure 2 shows plots which summarise the viscoelastic behaviour of the model for hexagonal grains. The top two panels show the scaled absolute shear modulus $|G^{*}|/G_{0}$ and the attenuation $Q^{-1}\equiv J_{2}/J_{1}$, quantities commonly plotted in experimental studies of forced oscillation; the bottom two panels show scaled versions of the compliances which illustrate the nature of the relaxation. Figure 2 also shows fits to the finite element results using some simple viscoelastic models: the Maxwell model, the Andrade model and the extended Burgers Model (see Appendix H for a description of these models).
The behaviour of the model is best understood by considering its extremes. As remarked earlier, the behaviour at zero frequency is that of a viscous fluid. The behaviour at low frequencies $\omega\tau_{M}\ll 1$ more generally (referred to as the diffusionally-accommodated regime by [14, 1]) is that of a Maxwell model with a dashpot of viscosity $\eta$ in series with a spring of compliance $J_{ss}$ or modulus $G_{ss}=1/J_{ss}$. $\eta$ is the steady-state viscosity and $G_{ss}$ is the steady-state modulus. The steady-state compliance $J_{ss}$ is the limiting value of the storage compliance: $J_{1}\rightarrow J_{ss}$ as $\omega\tau_{M}\rightarrow 0$. The relaxation strength $\Delta$ is the fractional increase in compliance due to the relaxation of stresses by grain boundary diffusion, given by
@@
At high frequencies (the diffusionally-assisted regime of [14, 1]), with $\omega\tau_{M}\gg 1$, the behaviour is that of an Andrade model, with power-law variation of attenuation as $Q^{-1}\propto\left(\omega\tau_{M}\right)^{-\alpha}$ for a particular exponent $\alpha$. The Andrade exponent $\alpha$ can be calculated explicitly using singularity analysis [15, 16, 17] following a scaling argument due to Lee et al. [7]. In the absence of grain boundary diffusion, the pure elastic problem with freely-sliding grain-boundaries has singularities in the stress at the triple junctions where three grains meet, i.e. the normal stresses go as $\sigma_{nn}\propto r^{-\lambda}$ where $r$ is the distance from the triple junction. The analysis of Picu and Gupta [15] shows that for a triple junction with equal angles of 120° and freely sliding grain boundaries the exponent $\lambda=0.4491862$.
In the presence of diffusion, the singularity in the stress is removed. The length scale over which diffusion acts depends on the frequency of the forcing and scales with (Appendix K)
@@
The other parameter which specifies the Andrade model is the Andrade time $\tau_{A}$ (appendix H.4). Unlike the exponent $\alpha$, this time cannot be determined purely by examining the singular behaviour at the triple junctions, but is also sensitive to the overall grain geometry and Poisson’s ratio. Figure 1c plots the Andrade time calculated from the finite element analysis as a function of Poisson’s ratio, and shows that the Andrade time is broadly similar to the Maxwell time, increasing from $\tau_{A}/\tau_{M}=0.36$ to 2.13. The dimensionless ratio $\tau_{A}/\tau_{M}$ is sometimes referred to as the Andrade $\zeta$ parameter in tidal heating studies [18]. For a typical rock value of $\nu=0.3$, $\tau_{A}/\tau_{M}=1.315$.
The Maxwell model and the Andrade model characterise the behaviour of the finite element results at low and high dimensionless frequency respectively. Also shown in Figure 2 is another model fit to the finite element data using an extended Burgers Model (appendix H.5, [19]). This model is similar to the Andrade model in having a relaxation spectrum with a power-law distribution of relaxation times, but has a long-period cut-off to the relaxation spectrum given by a parameter $\tau_{H}$ which can be related to $\tau_{A}$ and $\Delta$ by $\tau_{H}=\tau_{A}\left(\Gamma(1-\alpha)\Delta\right)^{1/\alpha}$, where $\Gamma(z)$ is the Gamma function. The extended Burgers model has the same asymptotic behaviour as the finite element model at both low and high frequencies. As can be seen in Figure 2, the extended Burgers model provides an excellent fit to the finite element results, with only subtle differences around $\omega\tau_{M}=20$ to be seen in panels a and b which show absolute shear modulus and attenuation. The differences are somewhat more noticeable in the plots of scaled compliance in panels c and d, but even in these plots the differences are slight. Thus an effective parametrisation of the finite element results is to use an extended Burgers model with the appropriate values of $G_{0}$, $\tau_{M}$, $\Delta$ (Figure 1b), $\alpha=0.3672092$ and $\tau_{A}/\tau_{M}$ (Figure 1c).
@@ 5 Discussion
A more significant difference between the present study and that of Lee et al. [7] is the values of $\tau_{A}/\tau_{M}$ which are significantly larger for the bicrystal models of [7] than the models here; that is to say the bicrystal models of [7] show significantly less attenuation for the same dimensionless frequency (the grey lines are lower than the black lines on Figure 6b). This difference must arise from the overall differences in grain geometry, where it should be noted that the bicrystal model has two lengthscales which define it rather than just one for the model here: there is one lengthscale associated with the variations of the interface, and another associated with the width of the bicrystal. Lee et al. [7] present results where the width is 5 times that of the scale of variations on the interface; it is expected that the ratio of these two lengthscales plays an important role in determining the amount of attenuation, as dissipation only takes place on the boundary but elastic energy is stored throughout the grains.
@@
Figure 7: Andrade exponent $\alpha$ as a function of junction angle $\theta$ for a triple junction with angles $\theta$, $180\text{\textdegree}-\theta/2$ and $180\text{\textdegree}-\theta/2$. Dotted lines indicate the junction angles for the tetrakaidecahedron ($\theta=109.47\text{\textdegree}$, $\alpha=\tfrac{1}{3}$) and the hexagon ($\theta=120\text{\textdegree}$, $\alpha=0.3672092$).
It is not clear why some laboratory studies show lower Andrade exponents than the models here. In the models, the Andrade exponent is purely controlled by the triple junction geometries, and significant departures from the symmetric 120°-120°-120° junction angles are needed to produce Andrade exponents of 0.2. This is illustrated in Figure 7, which plots the Andrade exponent expected for asymmetric junctions with one angle of $\theta$ and two identical angles of $180\text{\textdegree}-\theta/2$. An exponent of 0.2 could be produced by having a $60\text{\textdegree}-150\text{\textdegree}-150\text{\textdegree}$ junction, but this seems far from what is observed in the microstructures. There may be some other source of anisotropy or geometrical feature that influences the exponent in some experiments, or there may be another dissipation mechanism producing attenuation that is mapped by the fitting to produce a lower effective Andrade exponent.
@@
[…]
@@
Understanding the viscoelastic behaviour of rocks at high homologous temperatures is of particular relevance to Earth sciences when interpreting the significant variations in seismic wavespeed, as well as increased attenuation, that occur near the lithosphere-asthenosphere boundary [e.g. 31, 26]. The models here do not produce enough attenuation to explain what is seen in seismology. For example, with a steady-state asthenospheric shear viscosity $\eta=2\times 10^{19}$~Pa s and an intrinsic shear modulus $\mu=70$ GPa, the Maxwell time $\tau_{M}=3\times 10^{8}$ s, i.e. around 10 years. For a surface wave with a period around T=50T=50~s, the dimensionless frequency is $\omega\tau_{M}=2\pi\tau_{M}/T=4\times 10^{7}$, for which the models yield $Q^{-1}\sim 10^{-3}$ (Figure 6), significantly less than the asthenospheric $Q^{-1}\sim 10^{-2}$ estimated by Ma et al. [31] and given in PREM [32]. Previous work in extrapolating laboratory experiments by Jackson and Faul [19] suggested that diffusionally-accommodated grain boundary sliding can account for the $Q^{-1}\sim 10^{-2}$ level of attenuation at seismic frequencies. This extrapolation was based on a proposed variation of Maxwell time as linear with grain size, $\tau_{M}\propto d$, rather than $\tau_{M}\propto d^{3}$ as expected from theory for grain-boundary diffusion. More recent experiments in the same laboratory [33] have since observed a much stronger grain size dependence with $\tau_{M}\propto d^{3.17}$, closer to the behaviour predicted by theory, and for which extrapolation does not yield sufficient attenuation at seismic frequencies. Thus additional mechanisms of dissipation, such as those associated with premelting [22, 27], impurities [34], elastically-accommodated grain boundary sliding [35], or dislocations [36, 37] are needed to explain the seismic observations.
@@
There are a number of natural ways the present modelling work can be extended. One natural extension is to consider diffusion within the interior of grains rather than along the grain boundaries. An example of such a calculation for hexagons is given in Appendix L, and the most notable result is that the relationship between the Andrade exponent and the stress singularity exponent changes to $\alpha=1-\lambda$, such that $\alpha=1/2$ is expected for tetrakaidecahedral grains. A key flexibility of the finite element modelling approach taken here is that can be adapted to model a wide range of different grain shapes with potentially anisotropic properties. Calculations could be performed using a synthetic rock texture which has a range of grain sizes, e.g. using Neper [38]. The expectation would be that the Andrade exponent would remain determined by the triple junction angles, but that the effective $\tau_{A}/\tau_{M}$ and $\Delta$ parameters may change. All the calculations here are for materials in solid-state; solid both in the grain interiors and on the grain boundaries. The effect of melt and melt geometry [20] could be explored, as was done for the case of steady-state diffusion creep viscosity in [8]. The presence of melt allows for a steady-state bulk viscosity of the material, and so calculations with melt would allow for both shear and bulk attenuation phenomenon to be explored. However, perhaps most interesting would be to consider the temperature dependent properties of the grain boundaries and the triple lines and explore the effect of premelting and sub-solidus mechanisms of attenuation more generally.